# VLM-SFD: VLM-Assisted Siamese Flow Diffusion Framework for Dual-Arm Cooperative Manipulation

### v1

@@ I Introduction
Dual-arm cooperative manipulation involves the synchronized control of two robotic arms to execute complex tasks with high precision and adaptability [he2022double, kruger2011dual, bai2022anthropomorphic]. Compared to the extensively-studied field of single-arm manipulation [im2flow2act, vlmpc, chi2024universal], dual-arm manipulation poses significant challenges in coordination and control, requiring precise synchronization, collision avoidance, and adaptive planning [dastider2024apex].
Previous research has explored various directions, including control optimization [jinjun2019symmetrical, ren2024enabling, abbas2023systematic], vision-based learning [zhao2023learning, liu2024rdt], and advanced motion planning frameworks [sintov2020motion, yu2023coarse]. While these methods enhance adaptability and synchronization, they often rely on extensive dual-arm demonstrations and require meticulous tuning, limiting their generalization capabilities. More recently, flow-based policies [orion, im2flow2act] and diffusion models [im2flow2act] have shown promise in robotic manipulation, demonstrating improved motion prediction while requiring fewer human demonstrations. Despite these advancements, extending flow-based methods to dual-arm cooperative manipulation presents significant yet largely unaddressed research challenges including modeling the interactions between two manipulators, ensuring synchronized coordination while avoiding collisions, task allocations between manipulators based on spatial and temporal constraints, and lacking of sufficient real-world data of dual-arm demonstrations.
To overcome these challenges, we introduce the VLM-Assisted Siamese Flow Diffusion (VLM-SFD) framework. This framework enables the learning of coupled dual-arm behaviors from a small set of human demonstrations by leveraging the proposed Siamese Flow Diffusion Network (SFDNet) and a novel VLM-assisted spatial-temporal task allocation strategy. The SFDNet employs a dual-encoder-decoder architecture to generate object-centric motion flows through a conditional diffusion process. Initially, target objects are localized and encoded by a Siamese Variational Autoencoder (VAE) encoder, yielding latent features that capture object properties and interactions. These latent features are then fed into a Siamese UNet, conditioned on the task instruction, followed by a Siamese VAE decoder which iteratively refines the motion flows through diffusion, generating two synchronized and compatible motion streams that guides dual-arm coordination.
@@
We evaluated the VLM-SFD framework on four challenging real-world tasks, demonstrating its data efficiency and deployment advantages. Unlike prior imitation learning-based methods which often require extensive data and real-world fine-tuning, VLM-SFD achieves competitive performance with only ten human demonstrations for each task. Furthermore, our model can be directly deployed on a physical dual-arm robotic platform without any further fine-tuning. Experimental results confirm that VLM-SFD maintains high success rates and efficient execution, despite using significantly fewer training examples and requiring no extra real-world adaptation, underscoring its advantage in practical deployment scenarios.
@@
  • We introduce SFDNet, a dual-encoder-decoder Siamese Flow Diffusion Network that applies conditional diffusion for dual-arm action generation. To the best of our knowledge, this is the first framework to leverage diffusion-based motion synthesis in a dual-arm setting, achieving tightly synchronized motions from limited demonstration data.
@@ II-A Dual-Arm Cooperative Manipulation
Dual-arm cooperative manipulation involves two robotic arms working together through synchronized planning and task-space coordination for effective object manipulation. Previous work in this area has explored methods [jiao2022adaptive, zhao2023learning, yu2023coarse] from low-level control to high-level task execution. One major research direction focuses on the optimization of control strategies, including dynamical system modeling [ren2024enabling], hybrid control approaches [jiao2022adaptive, ren2017adaptive], and adaptive force interaction [lu2018fuzzy]. These methods enable precise force modulation and trajectory execution, making them particularly relevant for high-precision tasks such as robot-assisted surgery [wu2019coordinated], cloth manipulation [liu2024rdt], and assembly [wang2024dexcap]. While computer vision has advanced significantly in robotics, another emerging direction is vision-based learning, particularly through reinforcement learning and imitation learning, which making robots more intelligent and adaptable to complex tasks and environments [soori2023artificial]. Methods like DexCap [wang2024dexcap] and ALOHA [zhao2023learning] leverage teleoperation for efficient data collection of bimanual skills, while RDT-1B [liu2024rdt] and Bi-DexHands [chen2023bi] employ large-scale bimanual training datasets to improve dexterity across diverse scenarios. Although these methods offer improved scene adaptability and skill transfer, they often rely on extensive data collection to generalize effectively. Recently various motion planning frameworks have been developed for high-level task execution in dual-arm cooperative manipulation. Yu et al. [yu2023coarse] and Sun et al. [sun2022motion] both explored hierarchical motion planning to handle complex tasks with dynamic obstacle avoidance. While these methods improve adaptability and trajectory smoothness, they still struggle to generalize to unstructured environments and diverse tasks.
@@ III-A Overview: VLM-SFD Framework
Dual-arm cooperative manipulation requires two robotic arms to work synergistically on tasks involving multiple objects, enabling complex coorperative manipulation. To address this challenge, we propose a VLM-assisted Siamese flow diffusion framework, which formulates the problem as a two-stage process, as illustrated in Fig. 1. In the training stage, the system learns object-centric motion flows for the objects involved in the interaction. In the testing stage, the learned 2D motion flows are mapped into the 3D execution space of the robotic arms, ensuring that both arms move in precise coordination.
@@
where $F_{1}\in\mathbb{R}^{3\times T\times H\times W},F_{2}\in\mathbb{R}^{3\times T\times H\times W}$ is the extracted motion flows, $T$ denotes the temporal length of the video, and $H,W$ are the spatial dimensions of the extracted motion flow. Each motion flow encodes the per-pixel displacement of the target objects over time, offering a structured representation of their movement within the scene and serving as supervision for training the SFDNet.
In the testing stage, we deploy the trained SFDNet directly onto a real-world dual-arm robotic platform. The predicted 2D object-centric motion flows are first mapped into the 3D execution space. By leveraging VLM-assisted spatial-temporal knowledge reasoning, the framework dynamically interprets task semantics, object relationships, and execution dependencies to assign motion trajectories to each manipulator. This enables adaptive scheduling and collision-free coordination, optimizing the execution order based on both spatial constraints and temporal dependencies within the task.
The details of the proposed SFDNet and the VLM-assisted task assignment strategy are introduced in the following sections.
@@ III-B SFDNet: Siamese Flow Diffusion Network
$L_{\text{diffusion}}=\sum_{i}\mathbb{E}_{t,{\epsilon_{i}^{1:T}},\mathbf{z}_{i}^{1:T}(0)}\bigl{|}\bigl{|}{\epsilon}_{i}^{1:T}-\hat{\epsilon}_{i}^{1:T}\bigr{|}\bigr{|}^{2}.$ (7)
@@ III-C Inference with the VLM-Assisted Spatial-Temporal Task Allocation Strategy
During the testing stage, we deploy the trained SFDNet directly onto a real-world dual-arm platform. Following AnimateDiff [animatediff], we first inflate the initial observation frame $\mathcal{I}_{0}$ to match the temporal length of the video to obtain $\mathcal{I}_{0}^{1:T}$, then SFDNet generates the two-stream object-centric motion flows under the condition of task instruction $\mathcal{L}$ denoted as:
$\{\hat{F}_{1},\hat{F}_{2}\}=\mathrm{SFDNet}(\mathcal{I}_{0}^{1:T},\mathcal{L}),$ (9)
where $\hat{F}_{1},\hat{F}_{2}\in\mathbb{R}^{3\times T\times H\times W}$ are the predicted motion flows. Then we project the predicted 2D motion flows in to 3D execution space using depth information captured by the RGB-D camera, which can be formulated as:
$\mathrm{Traj}_{i}=\mathcal{G}(\mathbf{T}\hat{F}_{i}),\quad i\in\{1,2\},$ (10)
where $\mathbf{T}$ is the transformation matrix derived from the camera’s intrinsic and extrinsic parameters, mapping the 2D coordinates into 3D points, $\mathcal{G}(\cdot)$ represents the post-processing includes smoothing and merging operations, the final 3D trajectory $\mathrm{Traj_{i}}\in\mathbb{R}^{P\times 6}$ is the 3D trajectory containing $P$ waypoints defined by its position and orientation within 3D space.
@@
To address this, we introduce the VLM-Assisted Spatial-Temporal Task Allocation Strategy, leveraging a pre-trained VLM (GPT-4o [gpt4o]) for knowledge-driven reasoning. Specifically, instead of executing full-length trajectories simultaneously, each predicted trajectory is first segmented into smaller sub-trajectories for finer execution control. Then the VLM determines the optimal segmentation strategy by considering the workspace layout, object dependencies, and motion complexity. Given a predicted trajectory $\text{Traj}_{i}$, a pre-defined trajectory segmentation tool function can be invoked by VLM, which can be denoted as:
$\{\mathrm{Seg}_{i}^{1},\mathrm{Seg}_{i}^{2},...,\mathrm{Seg}_{i}^{M}\}=F_{\text{seg}}(\mathrm{Traj}_{i},M_{\text{VLM}}),$ (11)
where $M_{\text{VLM}}$ is the number of sub-trajectories VLM determines to dynamically suit the task, and each $\text{Seg}_{i}^{m}$ represents a sub-trajectory containing a fraction of the original trajectory’s waypoints.
Once trajectories are segmented, the VLM determines the optimal execution sequence by leveraging another tool function trajectory visualization that overlays the segmented sub-trajectories onto the initial frame using distinct colors and numerical labels to represent each segment. This allows VLM to directly analyze all sub-trajectories at once and determine an optimized execution order.
Since collision and cooperation constraints are inherently embedded in this ranking process, the prompted VLM ensures that (1) no two conflicting sub-trajectories are scheduled simultaneously; (2) dual-arm coordination is naturally optimized, preventing unnecessary delays; (3) execution is both efficient and collision-free without explicit constraint handling.
@@ IV-A Experimental Details
We collect training data by human demonstrations. Specifically, we adopt four challenging cooperative tasks following the baselines [zhao2023learning, chi2023diffusion]:
@@
Fig. 2 provides examples of each task. These tasks can be categorized into two types of dual-arm collaboration: synchronous (e.g., Packing) and asynchronous (e.g., Put X into Pot, Pouring, and Pulling Drawer & Placing). All manipulated objects are rigid or non-prehensile. For each task, we collect only 10 episodes of human demonstrations involving 3 distinct objects to be manipulated within a cluttered environment. The workspace contains numerous distractor items, requiring the robot to accurately identify the object specified by the language instruction while avoiding collisions during manipulation.
@@
Following Im2Flow2Act [im2flow2act], we modify and extend AnimateDiff [animatediff], an advanced text-to-image diffusion model, into the proposed siamese architecture and fine-tune it on our dataset. We adopt the AdamW [adamw] optimizer with a learning rate of 0.0001 and a weight decay of 0.01, training the model for a total of 4500 epochs. To enable efficient fine-tuning, we employ LoRA [lora] with a rank of 128. Input images are resized to $224\times 224$ before being fed into the model.
@@ IV-B Experiment Results
Furthermore, it is worth noting the advantage in training efficiency of the proposed VLM-SFD framework. Unlike ACT and DP, which require hundreds of binocular video recordings and robot trajectories per task, along with complex post-training like action-level decoding, our framework is only trained on a small set of easy-to-collect human demonstrations and can be directly deployed on the real-world robot. This comparison demonstrates the efficiency of our method in learning from limited data with low training cost.
TABLE I: Ablation Studies on Siamese Structure and Task Allocation
Method Pouring
@@ IV-C Ablation Study
We conduct ablation studies to access the contribution of key components in our VLM-SFD framework. The detailed results are summarized in Table I. Considering the need for fine-grained control and knowledge-driven spatial reasoning, we select the challenging pouring task for 20 repeat tests.

### v2

@@ I Introduction
Dual-arm cooperative manipulation involves the synchronized control of two robotic arms to execute complex tasks with high precision and adaptability [he2022double, kruger2011dual, bai2022anthropomorphic]. Compared to the extensively-studied field of single-arm manipulation [im2flow2act, chi2024universal], dual-arm manipulation poses significant challenges in coordination and control, requiring precise synchronization, collision avoidance, and adaptive planning [dastider2024apex].
Previous research has explored various directions, including control optimization [ibvs, tarbouriech2018dual], vision-based learning [zhao2023learning, liu2024rdt], and advanced motion planning frameworks [sintov2020motion, yu2023coarse]. While these methods enhance adaptability and synchronization, they often rely on extensive dual-arm demonstrations and require meticulous tuning, limiting their generalization capabilities. More recently, flow-based policies [orion, im2flow2act] and diffusion models [im2flow2act] have shown promise in robotic manipulation, demonstrating improved motion prediction while requiring fewer human demonstrations. Despite these advancements, extending flow-based methods to dual-arm cooperative manipulation presents significant yet largely unaddressed research challenges including modeling the interactions between two manipulators, ensuring synchronized coordination while avoiding collisions, task allocations between manipulators based on spatial and temporal constraints, and lacking of sufficient real-world data of dual-arm demonstrations.
To overcome these challenges, we introduce the VLM-Assisted Siamese Flow Diffusion (VLM-SFD) framework. This framework enables the learning of coupled dual-arm behaviors from a small set of human demonstrations by leveraging the proposed Siamese Flow Diffusion Network (SFDNet) and a novel VLM-assisted spatial-temporal task allocation strategy. The SFDNet employs a dual-encoder-decoder architecture to generate object-centric motion flows through a conditional diffusion process. Initially, target objects are localized and extracted motion flows which are then encoded by a Siamese Variational Autoencoder (VAE) encoder, yielding latent features that capture object properties and interactions. These latent features are then fed into a Siamese UNet, conditioned on the task instruction, followed by a Siamese VAE decoder which iteratively refines the motion flows through diffusion, generating two synchronized and compatible motion streams that guides dual-arm coordination.
@@
We evaluated the VLM-SFD framework on four challenging real-world tasks, demonstrating its data efficiency and deployment advantages. Unlike prior imitation learning-based methods which often require extensive data and real-world fine-tuning, VLM-SFD achieves competitive performance with only ten human demonstrations for each task. Furthermore, our model can be directly deployed on a physical dual-arm robotic platform without further fine-tuning. Experimental results confirm that VLM-SFD maintains high success rates and efficient execution, despite using significantly fewer training examples and requiring no extra real-world adaptation, underscoring its advantage in practical deployment scenarios.
@@
  • We introduce SFDNet, a dual-encoder-decoder Siamese Flow Diffusion Network that applies conditional diffusion for dual-arm action generation. To the best of our knowledge, this is the first framework to leverage diffusion-based motion synthesis in a dual-arm setting, enabling flexible coordination and synchronized or sequential behaviors from limited demonstration data.
@@ II-A Dual-Arm Cooperative Manipulation
Dual-arm cooperative manipulation involves two robotic arms working together through synchronized planning and task-space coordination for effective object manipulation. Previous work in this area has explored methods [jiao2022adaptive, zhao2023learning, yu2023coarse] from low-level control to high-level task execution. One major research direction focuses on the optimization of control strategies, including dynamical system modeling [ren2024enabling], hybrid control approaches [jiao2022adaptive, ren2017adaptive], and adaptive force interaction [lu2018fuzzy]. These methods enable precise force modulation and trajectory execution, making them particularly relevant for high-precision tasks such as robot-assisted surgery [wu2019coordinated], cloth manipulation [liu2024rdt], and assembly [wang2024dexcap]. While computer vision has advanced significantly in robotics, another emerging direction is vision-based learning, particularly through reinforcement learning and imitation learning, which making robots more intelligent and adaptable to complex tasks and environments [soori2023artificial]. Methods like DexCap [wang2024dexcap] and ALOHA [zhao2023learning] leverage teleoperation for efficient data collection of bimanual skills, while RDT-1B [liu2024rdt] and Bi-DexHands [chen2023bi] employ large-scale bimanual training datasets to improve dexterity across diverse scenarios. Although these methods offer improved scene adaptability and skill transfer, they often rely on extensive data collection to generalize effectively. In addition to learning-based approaches, traditional control methods like sparse kinematic control [tarbouriech2018dual] and IBVS [ibvs] have been explored for dual-arm cooperation. Sparse kinematic control minimizes the number of actuated joints for efficient coordination, while IBVS integrates visual feedback to adjust manipulator poses, reducing synchronization errors. However, these methods are often task-specific and struggle to generalize across diverse scenarios. Recently, various motion planning frameworks have been developed for high-level task execution in dual-arm cooperative manipulation. Yu et al. [yu2023coarse] and Sun et al. [sun2022motion] both explored hierarchical motion planning to handle complex tasks with dynamic obstacle avoidance. While these methods improve adaptability and smoothness, they still struggle to generalize to unstructured environments and diverse tasks.
@@ III-A Overview: VLM-SFD Framework
Dual-arm cooperative manipulation requires two robotic arms to work synergistically on tasks involving multiple objects, enabling complex cooperative manipulation. To address this challenge, we propose a VLM-assisted Siamese flow diffusion framework, which formulates the problem as a two-stage process, as illustrated in Fig. 1. In the training stage, the system learns object-centric motion flows for the objects involved in the interaction. In testing stage, the learned 2D motion flows are mapped into the 3D execution space, ensuring that both arms move in precise coordination.
@@
where $F_{1}\in\mathbb{R}^{3\times T\times H\times W},F_{2}\in\mathbb{R}^{3\times T\times H\times W}$ is the extracted motion flows, where the first two dimensions are 2D coordinates of each point within flow and the last dimension notes the visibility of the point, $T$ denotes the temporal length of the video, and $H,W$ are the spatial dimensions of the extracted motion flow. Each motion flow encodes the per-pixel displacement of the target objects over time, offering a structured representation of their movement within the scene and serving as supervision for training the SFDNet.
In the testing stage, we deploy the trained SFDNet directly onto a real-world dual-arm platform. The predicted 2D object-centric motion flows are first mapped into the 3D execution space. By leveraging VLM-assisted spatial-temporal knowledge reasoning, the framework dynamically interprets task semantics, object relationships, and execution dependencies to assign motion trajectories to each manipulator. This enables adaptive scheduling and collision-free coordination, optimizing the execution order based on spatial constraints and temporal dependencies within the task.
The details of the SFDNet and the VLM-assisted task assignment strategy are introduced in the following sections.
@@ III-B SFDNet: Siamese Flow Diffusion Network
[…]
@@ III-C Inference with the VLM-Assisted Spatial-Temporal Task Allocation Strategy
During the testing stage, we deploy the trained SFDNet directly onto a real-world dual-arm platform. Following AnimateDiff [animatediff] and Im2Flow2Act [im2flow2act], we first randomly initialized the latent code sequences $\mathbf{z}^{1:T}_{1}(0)$ and $\mathbf{z}^{1:T}_{2}(0)$ from the standard normal distribution, then SFDNet generates the two-stream object-centric motion flows under the condition of task instruction $\mathcal{L}$ denoted as:
$\{\hat{F}_{1},\hat{F}_{2}\}=\mathrm{SFDNet}(\mathbf{z}^{1:T}_{1}(0),\mathbf{z}^{1:T}_{2}(0),\mathcal{L}),$ (9)
where $\hat{F}_{1},\hat{F}_{2}\in\mathbb{R}^{3\times T\times H\times W}$ are the predicted motion flows. We then follow AVDC [ko2023learning] to reconstruct 3D trajectories. We first transform 2D points from motion flows in the initial frame into 3D points using the camera’s intrinsic parameters and depth information captured by the RGB-D camera. For each subsequent frame, we calculate the following trajectory waypoints by minimizing the reprojection error with randomly initialized transformation matrix $\mathbf{T}_{t}\in\mathbb{R}^{4\times 4}$ that describes the transformation of 3D waypoint from the first frame to $t^{\text{th}}$ frame. The reprojection error is defined as:
$E_{t}=\sum_{i=1}^{N}\left\|\mathbf{p}_{i,t}-\Pi(\mathbf{T}_{t}\cdot\mathbf{P}_{i,0})\right\|^{2},$ (10)
where $N$ is the number of points within each frame of the 2D flow, $\mathbf{p}_{i,t}=(x_{i,t},y_{i,t})$ is $i$-th point in frame $t$, $\Pi(\cdot)$ is the projection function mapping 3D points back to the 2D image plane, $\mathbf{P}_{i,0}$ represents the 3D coordinates of the $i$-th point in the first frame. The term $E_{t}$ represents the reprojection error for frame $t$. The optimized $\{\mathbf{T}_{0},...,\mathbf{T}_{T}\}$ can further derive 3D trajectories $\mathrm{Traj}_{i}\in\mathbb{R}^{T\times 6}$ for two objects ($i\in\{1,2\}$) that describes $T$ 3D waypoints with coordinates as well as rotation angle. These trajectories are refined through post-processing operations, such as smoothing and merging, to ensure coherence.
@@
To address this, we introduce the VLM-Assisted Spatial-Temporal Task Allocation Strategy, leveraging a pre-trained VLM (GPT-4o [gpt4o]) for knowledge-driven reasoning. Specifically, instead of executing full-length trajectories simultaneously, we first overlay the predicted 2D flows on the initial frame $\mathcal{I}_{0}$. Based on this visualization, the VLM determines the allocation of flows to each arm and decides the segmentation granularity $M_{\text{VLM}}$, and then calls a pre-designed trajectory segmentation tool $F_{\text{seg}}$ to partition each predicted flow into smaller 2D segments:
$\{\mathrm{Seg}_{i}^{1},\mathrm{Seg}_{i}^{2},...,\mathrm{Seg}_{i}^{M_{\text{VLM}}}\}=F_{\text{seg}}(\hat{F}_{i},M_{\text{VLM}}),$ (11)
where each $\mathrm{Seg}_{i}^{m}$ represents a fraction of the original flow. The segments are visualized on $\mathcal{I}_{0}$ with distinct colors and numeric labels, allowing the VLM to jointly analyze all segments and rank them for an optimized execution order, then assigning the open/close control signal for the grippers at the start or end of specific segments. Since each waypoint in the 2D flow corresponds one-to-one with a 3D trajectory waypoint, the selected sub-trajectories can be directly mapped back into 3D space for execution.
Since collision and cooperation constraints are inherent to the ranking process, the proposed strategy ensures that (1) no two conflicting sub-trajectories are scheduled simultaneously; (2) dual-arm coordination is naturally optimized, preventing unnecessary delays; (3) execution is both efficient and collision-free without explicit constraint handling.
@@ IV-A Experimental Details
We collect training data by human demonstrations. Specifically, we adopt four challenging tasks following the baselines [zhao2023learning, chi2023diffusion]:
@@
Fig. 2 provides examples of each task. These tasks can be categorized into two types of dual-arm collaboration: synchronous (e.g., Packing) and asynchronous (e.g., Put X into Pot, Pouring, and Pulling Drawer & Placing). All manipulated objects are rigid or minimally deformable. For each task, we collect only 10 episodes of human demonstrations involving 3 distinct objects to be manipulated within a cluttered environment. On average, the demonstration lengths are about 120 frames ($\approx$ 4.8 seconds at 25 FPS). The workspace contains numerous distractor items, requiring the robot to accurately identify the object specified by the language instruction while avoiding collisions during manipulation.
@@
Following Im2Flow2Act [im2flow2act], we modify and extend AnimateDiff [animatediff], an advanced text-to-image diffusion model, into the proposed siamese architecture and fine-tune it on our dataset.
We adopt the AdamW [adamw] optimizer with a learning rate of 0.0001 and a weight decay of 0.01. To enable efficient fine-tuning, we employ LoRA [lora] with a rank of 128. Input images are resized to $224\times 224$ before being fed into the model. The training was conducted on a server node equipped with two Intel Xeon Gold 6326 processors and two NVIDIA A100 40GB GPUs. The total training time was approximately 60 GPU hours on this setup.
@@ IV-B Experiment Results
Furthermore, it is worth noting the advantage in training efficiency of the proposed VLM-SFD framework. Unlike ACT and DP, which require hundreds of binocular video recordings and robot trajectories per task, along with complex post-training like action-level decoding, our framework is trained on a small set of easy-to-collect human demonstrations and can be directly deployed on the real-world robot. This comparison demonstrates the efficiency of our method in learning from limited data with low training cost.
To investigate the time cost of the proposed method, we measured the average planning time for flow generation, VLM inference, and the execution time, as shown in Table I. These results demonstrate that despite the latency introduced during the planning phase due to the use of the pretrained VLM, the system can still support real-time task execution.
TABLE I: Planning and Execution Times for Four Tasks
Task Planning Time (s) Execution Time (s)
Put X into Pot 128.4 54.3
Packing 141.9 24.8
Pouring 176.3 96.4
Pulling Drawer & Placing 146.6 64.3
TABLE II: Ablation Studies on Siamese Structure and Task Allocation
Method Success Rate (%)
@@ IV-C Ablation Study
We conduct ablation studies to access the contribution of key components in our VLM-SFD framework. The detailed results are summarized in Table II. Considering the need for fine-grained control and knowledge-driven spatial reasoning, we select the challenging pouring task for 20 repeat tests.
@@
To further examine the influence of the number of demonstrations on task execution, we conducted ablation experiments on the Pouring and Put X into Pot tasks. As shown in Table III, for Pouring, the success rate increases from 40% with 5 demos to 65% with 10 demos, remaining stable with 20 demos. For the Put X into Pot task, performance similarly improves from 20% with 5 demos to 55% with 10 demos, and with only a marginal gain to 60% (just one additional successful trial) at 20 demos. These results suggest that around 10 demonstrations are sufficient to achieve stable task execution while balancing training cost and performance.
TABLE III: Ablation on number of demonstrations
Task 5 demos 10 demos 20 demos
Pouring 40% (8/20) 65% (13/20) 65% (13/20)
Put X into Pot 20% (4/20) 55% (11/20) 60% (12/20)